Is the 2-cycle (1 2) in the subgroup H=, T> 0 9 generated by \sigma and \tau? Explain answer.

(123)(45)(6789)
(13)(578)(49)
Answer #1

=(123)(45) (6789) e = (13) (578) (49) or=:(123) (45) (6789] (13) (578) (49) =(12) (475968) order af 6 7) = eam af order of (1